OpenStudy (love_333):

Find all the zeroes of the equation: -3x^4 + 27x^2 + 1200 = 0

OpenStudy (sshayer):

\[-3\left( x^4-9x^2-400 \right)=0\] \[x^2=t,x^4=t^2\] \[t^2-9t-400=0\] 1*-400=-400 25-16=9 25*-16=-400 \[t^2-\left( 25-16 \right)t-400=0\] \[t^2-25t+16t-400=0\] t(t-25)+16(t-25)=0 \[\left( t-25 \right)\left( t+16 \right)=0\] t=-16,25 \[x^2=-16=16 \times -1=16 \iota^2,x^2-16 \iota^2=0\] \[\left( x-4\iota \right)\left( x+4 \iota \right)=0,x=\pm \iota \] \[x^2-25=0,\left( x+5 \right)\left( x-5 \right)=0,x=\pm 5\]

OpenStudy (sshayer):

correction it is \[x= \pm 4 \iota ,\] 4 is missing.

OpenStudy (love_333):

@sshayer sorry, I'm not sure I understand, whats the answer?

OpenStudy (sshayer):

\[zeroes \ are \ x = \pm 5, \pm 4 \iota \]

OpenStudy (love_333):

wait, I'm sorry whats the thing at the end? It looks like an i without a dot?

OpenStudy (love_333):

Sorry, this is all new to me and I don't understand any of it

OpenStudy (love_333):

@marcelie @Kainui Do you guys have any ideas?

OpenStudy (sunnnystrong):

@Love_333 ... i is a complex number

OpenStudy (sshayer):

\[\iota=\sqrt{-1}\]